write an equation that requires the use of the addition property of equality, where 6 must be added to each side to solve the equation, and the solution is a negative number. the book say's x-6= -8..please explain the process.
Answer by ankor@dixie-net.com(22740)

write an equation that requires the use of the addition property of equality, where 6 must be added to each side to solve the equation, and the solution is a negative number. the book say's x-6= -8..please explain the process.
:
x - 6 = -8
:
In order to get x by itself on the left of the = sign we add 6 to both sides:
x - 6 + 6 = - 8 + 6
:
That eliminates the 6, do the math on the right and you have:
x = -2
:
The "addition property of equality," just means that whatever you add to
one side of an equation you add the same to the other side, thus preserving
equality, and in the process, you have isolated the variable (x) by itself,
thus finding x's value, which in most case "solves" the equation